# Beyond-mean-field phases of rotating dipolar condensates

### v1

@@ IV.4 $N=5$ bosons
Figure 13: The different phases obtained from the simulation of $N=5$ dipolar-interacting bosons in a rotating potential as a function of interaction strength $g_{d}$ and rotation speed $\Omega$ (on a logarithmic scale). The different panels show results obtained for different number of orbitals: (a) $M=1$ orbitals, (b) $M=2$ orbitals, (c) $M=5$ orbitals. (d) $M=10$ orbitals.
@@
However, these mean-field solutions are erroneous. Only by sizably increasing the number of orbitals to $M=5$ [Fig. 13(c)] and $M=10$ [Fig. 13(d)] do the true ground states appear in the regimes of fast rotations and strong interactions. These states consist mainly of exotic fragmented superfluids and cluster states. The fragmented superfluid take “blood cell” or droplet forms, as indicated in Fig. 14(b),(c). We remark that these shapes were already found in other studies with a larger particle number [93]. Here, we discover that they persist for systems with as little as five bosons. The cluster states appear in two different configurations with five-fold [Fig. 14(f)] or four-fold [Fig. 14(g)] symmetry. We did not find a systematic criterion for when one or the other symmetry is preferred in our simulations, and thus suspect that these states are quasidegenerate.
@@ V Conclusions and Outlook
Our work underscores the necessity of beyond-mean-field methods in capturing the intricate physics of rotating dipolar condensates and sets the stage for further theoretical and experimental explorations of these unconventional quantum phases in mesoscopic ultracold atomic systems. One immediate research direction to pursue is a detailed analysis of the correlation patterns in the identified phases, particularly focusing on the interplay between long-range interactions and vortex clustering [97, 98, 35, 99, 34, 100]. A key open question is whether the cluster states observed in this study exhibit true supersolid ordering, characterized by simultaneous phase coherence and density modulation [18, 22, 20, 19, 30, 43, 42]. Further MCTDH investigations, or the use of other techniques such as quantum Monte Carlo [101, 102, 103], could help confirm the nature of these phases and their potential experimental realizability.
@@
Another important aspect to consider is the exploration of how the full angular dependence of dipole-dipole interactions affects the stability and structure of the different phases [4, 104, 103, 16, 105, 35]. In our study, we considered a simplified model where interactions are effectively isotropic and repulsive in the plane, which corresponds to dipoles placed perpendicularly to it. A more complete treatment including anisotropic dipolar couplings could yield even richer phenomenology. Additionally, studying the dynamical response of the different states we encountered to time-dependent protocols – such as sudden quenches or periodic drives – could provide insights into the nonequilibrium behavior of dipolar gases. A particularly interesting direction involves continuous slowdowns, mimicking astrophysical processes like pulsar glitches. Given that recent studies have linked pulsar glitches to rotating supersolids [45, 46], it is compelling to ask whether cluster states combined with vortices exhibit qualitatively distinct dynamical signatures in decelerating condensates.
@@
Beyond bosonic systems, it would be fascinating to explore the fate of similar structures in fermionic gases subjected to rotation, particularly in regimes where the rotational speed approaches the Fermi velocity [106, 107, 108, 109, 104, 110]. This could give rise to unconventional quantum Hall-like states or exotic topological phases stabilized by dipolar interactions [41, 111, 112].
@@
Lastly, an exciting prospect is extending our study to cavity-mediated long-range interactions, as encountered in condensates coupled to optical cavities [113, 114, 115]. This setting introduces competition between vortex physics and superradiance, a problem that has so far been mostly addressed within mean-field approaches [116]. Understanding the impact of quantum fluctuations in such cavity-QED setups could reveal novel phases with strong light-matter coupling effects, offering connections between dipolar quantum gases and emergent photonic materials.
@@ Appendix A MCTDH-X
In this appendix, we provide a summary of the numerical method used to compute the ground states for the few-boson systems discussed in the main text. We employ the MultiConfigurational Time-Dependent Hartree (MCTDH) method for indistinguishable particles, coded in the MCTDH-X software [78, 80, 81, 82, 83, 84, 85]. This approach is tailored to solve the many-body Schrödinger equation and is particularly suited for investigating the ground state properties and dynamics of interacting ultracold systems, including dipolar gases. To date, MCTDH-X has been successfully applied to a range of ultracold atomic systems of various kind, from noninteracting Pauli crystals [117], to short-range interacting systems [118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 44, 130, 131], to dipolar interacting atoms and molecules [132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147], to ultracold gases coupled with cavity fields [148, 149, 150, 151, 152, 153, 154, 155, 156]. In this appendix, we will explicitly focus on bosonic systems only (MCTDHB), although MCTDH-X has been applied to fermionic gases and spinor condensates too.
@@
The number of orbitals governs both the accuracy and the convergence of the MCTDH expansion. For $M=1$, the ansatz reduces to a mean-field Gross-Pitaevskii equation (GPE) with a single (global) wave function. For $M>1$, many-body correlations can be captured effectively beyond mean-field approximations. Strictly speaking, the method is exact when $M\to\infty$. However, very often it is possible to obtain numerically exact results with a finite (and often small) number of $M$ when the many-body system is well-contained in the variational subspace spanned by the orbitals [79, 81]. The number of permanents in the MCTDHB decomposition (i.e. the different configurations in our variational subspace) is what determines the computational complexity of the problem, and it scales combinatorially with the particle and orbital number: $\left(\begin{array}[]{c}N+M-1\\ N\end{array}\right)$. This scaling is very unfavorable when either $M$ or $N$ (or both) are very large, and therefore acts as a computational limit for the applicability of the method to arbitrary systems. In practice, $M$ is chosen to balance precision requirements with computational feasibility.
To obtain the many-body state from the MCTDH decomposition, the Dirac-Frenkel time-dependent variational principle is applied to the Schrödinger equation written with the ansatz for the given many-body Hamiltonian [157]. This leads to a set of coupled integro-differential equations in the variational parameters, namely both the expansion coefficients $C_{\mathbf{n}}(t)$ and the orbitals $\psi_{i}(x;t)$. Solving these differential equations with optimized methods in real time allows us to compute the time evolution of a given initial state. If they are solved in imaginary time instead, we can obtain a variational approximation to the ground state. In this work, we focus on the latter approach.
@@
$\rho_{kq}=\begin{cases}\sum_{\mathbf{n}}|C_{\mathbf{n}}|^{2}n_{k},\quad&k=q,\\ \sum_{\mathbf{n}}C_{\mathbf{n}}^{*}C_{\mathbf{n}^{k}_{q}}\sqrt{n_{k}(n_{q}+1)},\quad&k\neq q,\\ \end{cases}$ (19)

### v2

@@ IV.4 $N=5$ bosons
Figure 13: The different phases obtained from the simulation of $N=5$ dipolar-interacting bosons in a rotating potential as a function of interaction strength $g_{d}$ and rotation speed $\Omega$ (on a logarithmic scale). The different panels show results obtained for different number of orbitals: (a) $M=1$ orbitals, (b) $M=2$ orbitals, (c) $M=5$ orbitals. (d) $M=10$ orbitals. The red letters in the phase diagram refer to the points in parameter space where the density distributions of Fig. 14 were probed.
@@
However, these mean-field solutions are erroneous. Only by sizably increasing the number of orbitals to $M=5$ [Fig. 13(c)] and $M=10$ [Fig. 13(d)] do the true ground states appear in the regimes of fast rotations and strong interactions. These states consist mainly of exotic fragmented superfluids and cluster states. The fragmented superfluid take “blood cell” or droplet forms, as indicated in Fig. 14(b),(c). We remark that these shapes were already found in other studies with a larger particle number [Schmidt:2021]. Here, we discover that they persist for systems with as little as five bosons. The cluster states appear in two different configurations with five-fold [Fig. 14(f)] or four-fold [Fig. 14(g)] symmetry. We did not find a systematic criterion for when one or the other symmetry is preferred in our simulations, and thus suspect that these states are quasidegenerate.
@@ V Conclusions and Outlook
Our work underscores the necessity of beyond-mean-field methods in capturing the intricate physics of rotating dipolar condensates and sets the stage for further theoretical and experimental explorations of these unconventional quantum phases in mesoscopic ultracold atomic systems. One immediate research direction to pursue is a detailed analysis of the correlation patterns in the identified phases, particularly focusing on the interplay between long-range interactions and vortex clustering [Prasad:2019, Klaus:2022, Bland:2023, Prasad:2024, Sabari:2024, Villois:2024]. A key open question is whether the cluster states observed in this study exhibit true supersolid ordering, characterized by simultaneous phase coherence and density modulation [Tanzi:2019, Tanzi:2019-2, Guo:2019, Chomaz:2019, Boettcher:2019, Poli:2024, Casotti:2024]. Further MCTDH investigations, or the use of other techniques such as quantum Monte Carlo [Pilati:2005, Buchler:2007, Macia:2014], could help confirm the nature of these phases and their potential experimental realizability.
@@
Another important aspect to consider is the exploration of how the full angular dependence of dipole-dipole interactions affects the stability and structure of the different phases [Lahaye:2009, Baranov:2012, Macia:2014, Lu:2015, Prasad:2021, Bland:2023]. In our study, we considered a simplified model where interactions are effectively isotropic and repulsive in the plane, which corresponds to dipoles placed perpendicularly to it. A more complete treatment including anisotropic dipolar couplings could yield even richer phenomenology. Additionally, studying the dynamical response of the different states we encountered to time-dependent protocols – such as sudden quenches or periodic drives – could provide insights into the nonequilibrium behavior of dipolar gases. A particularly interesting direction involves continuous slowdowns, mimicking astrophysical processes like pulsar glitches. Given that recent studies have linked pulsar glitches to rotating supersolids [Poli:2023, Bland:2024], it is compelling to ask whether cluster states combined with vortices exhibit qualitatively distinct dynamical signatures in decelerating condensates.
@@
Beyond bosonic systems, it would be fascinating to explore the fate of similar structures in fermionic gases subjected to rotation, particularly in regimes where the rotational speed approaches the Fermi velocity [Feder:2004, Zwierlein:2005, Tonini:2006, Cooper:2009, Baranov:2012, Kopyncinski:2021]. This could give rise to unconventional quantum Hall-like states or exotic topological phases stabilized by dipolar interactions [Mukherjee:2022, Predin:2023, Jiang:2024].
@@
Lastly, an exciting prospect is extending our study to cavity-mediated long-range interactions, as encountered in condensates coupled to optical cavities [Ritsch:2013, Mivehvar:2021, Schlawin:2022]. This setting introduces competition between vortex physics and superradiance, a problem that has so far been mostly addressed within mean-field approaches [Masalaeva:2024]. Understanding the impact of quantum fluctuations in such cavity-QED setups could reveal novel phases with strong light-matter coupling effects, offering connections between dipolar quantum gases and emergent photonic materials.
@@ Appendix A MCTDH-X
In this appendix, we provide a summary of the numerical method used to compute the ground states for the few-boson systems discussed in the main text. We employ the MultiConfigurational Time-Dependent Hartree (MCTDH) method for indistinguishable particles, coded in the MCTDH-X software [Alon:2008, Lode:2016, Fasshauer:2016, Lin:2020, Lode:2020, Molignini:2025-SciPost, MCTDHX]. This approach is tailored to solve the many-body Schrödinger equation and is particularly suited for investigating the ground state properties and dynamics of interacting ultracold systems, including dipolar gases. The power of MCTDH-X is underscored by the fact that it has been successfully applied to a broad range of ultracold atomic systems from noninteracting Pauli crystals [Xiang:2023], to short-range interacting systems [Roy:2018, Dutta:2019, Schaefer:2020, Lode:2021, Lode:2021-10, Debnath:2023, Roy:2023, Dutta:2023, Dutta:2024, Haldar:2024, Roy:2024-09, Roy:2024-11, Chatterjee:2024, Bhowmik:2025, Chakrabarti:2025-2], to dipolar interacting atoms and molecules [Fischer:2015, Chatterjee:2018, Chatterjee:2019, Bera:2019, Bera:2019-symm, Chatterjee:2020, Roy:2022, Hughes:2023, Bilinskaya:2024, Molignini:2024, Molignini:2024-2, Roy:2024-annals, Roy:2024-epjp, Molignini:2025-quasicryst1, Molignini:2025-quasicryst2, Chakrabarti:2025], to ultracold gases coupled with cavity fields [Lode:2017, Lode:2018, Molignini:2018, Lin:2019, Lin2:2019, Lin:2020-PRA, Lin:2021, Molignini:2022, Rosa-Medina:2022]. In this appendix, we will explicitly focus on bosonic systems only (MCTDHB), although MCTDH-X has been applied to fermionic gases and spinor condensates too.
@@
The number of orbitals governs both the accuracy and the convergence of the MCTDH expansion. For $M=1$, the ansatz reduces to a mean-field Gross-Pitaevskii equation (GPE) with a single (global) wave function. For $M>1$, many-body correlations can be captured effectively beyond mean-field approximations. Strictly speaking, the method is exact when $M\to\infty$. However, very often it is possible to obtain numerically exact results with a finite (and often small) number of $M$ when the many-body system is well-contained in the variational subspace spanned by the orbitals [Lode:2012, Fasshauer:2016]. The number of permanents in the MCTDHB decomposition (i.e. the different configurations in our variational subspace) is what determines the computational complexity of the problem, and it scales combinatorially with the particle and orbital number: $\left(\begin{array}[]{c}N+M-1\\ N\end{array}\right)$. This scaling is very unfavorable when either $M$ or $N$ (or both) are very large, and therefore acts as a computational limit for the applicability of the method to arbitrary systems. In practice, $M$ is chosen to balance precision requirements with computational feasibility.
To obtain the many-body state from the MCTDH decomposition, the Dirac-Frenkel time-dependent variational principle is applied to the Schrödinger equation written with the ansatz for the given many-body Hamiltonian [TDVM81]. This leads to a set of coupled integro-differential equations in the variational parameters, namely both the expansion coefficients $C_{\mathbf{n}}(t)$ and the orbitals $\psi_{i}(x;t)$. Solving these differential equations with optimized methods in real time allows us to compute the time evolution of a given initial state. If they are solved in imaginary time instead, we can obtain a variational approximation to the ground state. In this work, we focus on the latter approach.
@@
[…]
@@ Appendix B Units and system parameters
Table  1 summarizes the parameter values used to generate the density plots in Figs. 2, 6, 10, 14.
interaction $g_{d}[\bar{E}]$ velocity $\Omega[\bar{v}]$ orbitals $M$
$N=50$ (Fig. 2)
(a) 1.0 0.3 1
(b) 2.0 0.0 5
(c) 0.5 0.3 5
(d) 2.0 0.3 5
(e) 2.0 0.7 1
(f) 0.5 0.99 5
$N=20$ (Fig. 6)
(a) 0.1 0.5 1
(b) 5.0 0.0 5
(c) 2.0 0.1 5
(d) 1.0 0.3 5
(e) 0.1 0.99 5
(f) 5.0 0.99 5
$N=10$ (Fig. 10)
(a) 0.1 0.1 10
(b) 0.5 0.3 10
(c) 0.1 0.8 10
(d) 0.1 0.9 10
(e) 5.0 0.9 10
(f) 5.0 0.99 5
(g) 0.5 0.8 10
(h) 5.0 1.0 10
$N=5$ (Fig. 14)
(a) 0.5 0.5 2
(b) 0.5 0.95 10
(c) 0.1 0.95 10
(d) 1.0 0.7 10
(e) 2.0 0.95 1
(f) 2.0 0.95 10
(g) 5.0 0.3 10
(h) 5.0 1.0 10
Table 1: Parameters used to generate the density plots of Figs. 2, 6, 10, 14 in the main text.